OpenStudy (anonymous):

What effect does inflation have on your retirement income? at an inflation rate of 3% how much purchasing power would a 40,000 fixed income have in ten years?

OpenStudy (irishboy123):

discount the income by inflation after 1 yr PP = 40,000*(1-.03) after 2 yrs PP = 40,000*(1-.03)*(1-.03) after 2 yrs PP = 40,000*(1-.03)*(1-.03)*(1-.03)